An electric point charge of Q_1 = 9.63 nC is place
Answer #1

2.25 m Q1 Q2 Q3 x m

How, Q1 and Q2 are positive, the charge Q3 is negative then this must be between these two charges for that the net force to be zero.

F_{13}=k\frac{q_{1}q_{3}}{x^{2}}

F_{23}=k\frac{q_{2}q_{3}}{(2.25-x)^{2}}

k\frac{q_{1}q_{3}}{x^{2}}=k\frac{q_{2}q_{3}}{(2.25-x)^{2}}\rightarrow \frac{q_{1}}{x^{2}}=\frac{q_{2}}{(5.06-4.5x+x^{2})}

q_{1}(5.06-4.5x+x^{2})=q_{2}x^{2}\rightarrow

(9.63\times 10^{-9}C)(5.06-4.5x+x^{2})=(1.83\times 10^{-9}C)x^{2}

(9.63)(5.06-4.5x+x^{2})=(1.83)x^{2}\rightarrow 48.72-43.33x+9.63x^{2}=1.83x^{2}

48.72-43.33x+7.8x^{2}=0 Now solve the quadratic equation \frac{-b\pm \sqrt{b^{2}-4ac}}{2a}

\frac{-(-43.33)\pm \sqrt{(-43.33)^{2}-4(7.8)(48.72)}}{2(7.8)}

\frac{43.33\pm \sqrt{357.42}}{15.6}\rightarrow

\frac{43.33+\sqrt{357.42}}{15.6}\rightarrow x=3.98

\frac{43.33-\sqrt{357.42}}{15.6}\rightarrow x=1.56

The correct option is 1.56 m then the charge Q3 must be at 1.56 m of Q1 and 0.68 m from Q2

How, Q2 and Q3 are negative, the charge Q3 must be outside these two charges, but more close of the Q2 for that the net force to be zero.

F_{13}=k\frac{q_{1}q_{3}}{(2.25+x)^{2}}

F_{23}=k\frac{q_{2}q_{3}}{x^{2}}

k\frac{q_{2}q_{3}}{x^{2}}=k\frac{q_{1}q_{3}}{(2.25+x)^{2}}\rightarrow \frac{q_{2}}{x^{2}}=\frac{q_{1}}{(5.06+4.5x+x^{2})}

q_{2}(5.06+4.5x+x^{2})=q_{1}x^{2}\rightarrow

(1.83\times 10^{-9}C)(5.06+4.5x+x^{2})=(9.63\times 10^{-9}C)x^{2}

(1.83)(5.06+4.5x+x^{2})=(9.63)x^{2}\rightarrow 9.25+8.23x+1.83x^{2}=9.63x^{2}

9.25+8.23x-7.8x^{2}=0 Now solve the quadratic equation \frac{-b\pm \sqrt{b^{2}-4ac}}{2a}

\frac{-(8.23)\pm \sqrt{(8.23)^{2}-4(-7.8)(9.25)}}{2(-7.8)}

\frac{-8.23\pm \sqrt{356.33}}{-15.6}\rightarrow

\frac{-8.23+\sqrt{356.33}}{-15.6}\rightarrow x=-0.68m

\frac{-8.23-\sqrt{356.33}}{-15.6}\rightarrow x=1.73m

The correct option is 1.73 m then the charge Q3 must be at 1.73 m of Q2 and 3.98 m from Q1

If you have any question please let me know in the comments